OpenStudy (anonymous):

Need Help Understanding Vectors and Parametric Equations Are these parametric equations equal? x=2t-1 and y=-2t+2 VS x=-1-2t y=2+2t

OpenStudy (anonymous):

I had to figure out parametric and vector equations The book and I have different answers, but I think thats allowed with parametric equations.

OpenStudy (anonymous):

A(-1,2) and B(-3,4) are the two original points btw

OpenStudy (anonymous):

to figure out AB i figured it doesnt amtter if I do A-B or B-A

OpenStudy (amistre64):

a vector equations: given a direction vector <d> and an anchor point p L = p + <d>t

OpenStudy (amistre64):

vectors like to have positive first parts .... but yeah, it doesnt really matter overall

OpenStudy (anonymous):

But when I graphics those two sets of equations seperately They turned out to eb perpendicular to one another

OpenStudy (amistre64):

a parameteric setup takes the point and vector apart: x = px + dx*t y = py + dy*t z = pz + dz*t etc ...

OpenStudy (amistre64):

A(-1,2) and B(-3,4) +1-2 +1-2 -------- -2,2 ; or vector (1,-1)

OpenStudy (anonymous):

I'm just confused if I am getting some step wrong or if I am doing everything correctly, so I just have to check if x=2t-1 and y=-2t+2 is the same as x=-1-2t y=2+2t

OpenStudy (amistre64):

use either point, say p = (-1,2) L1 = (-1,2) + (1,-1)t x = -1+ t y = 2 - t

OpenStudy (anonymous):

The vector equation I got was (-1,2)+t(2,-2) While the book got (-1,2)+t(-2,2)

OpenStudy (anonymous):

Right I got that, but is my answer the same as the books in essence?

OpenStudy (anonymous):

Am I technically correct?

OpenStudy (anonymous):

Is (-1,2)+t(-2,2) the same as (-1,2)+t(-2,2)

OpenStudy (amistre64):

is (-1,2)+t(-2,2) the same as (-1,2)+t(-2,2) typo ??? they are the same

OpenStudy (anonymous):

no sorry Is (-1,2)+t(-2,2) the same as (-1,2)+t(2,-2)

OpenStudy (amistre64):

lol, they are the same, the vector the book gives is an exact result of A to B is all

OpenStudy (amistre64):

the direction vector need not be the same length since for any real value t, it gets scaled to reach any point in the line

OpenStudy (anonymous):

ok so it is correct regardless if it is 2,-2 or -2,2

OpenStudy (amistre64):

correct since: (2,-2) scaled by -1 gets us: (-2,2) they are the same "vector" as far as a line is concerned. If your book wants you to be more anal .... id suggest the books results

OpenStudy (anonymous):

but a vector is directional right>?

OpenStudy (anonymous):

no the book sint too anal

OpenStudy (amistre64):

yes, but when defining a line its immaterial ....

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

ok lol awesome. cuz my parametrics are all perpendicular and i was confused

OpenStudy (anonymous):

I thought if its a vector it always eneds direction

OpenStudy (amistre64):